OpenStudy (brooke2353):

3x^2 +10x + 8 pLEAASE HELP I WILL GIVE MEDAL

OpenStudy (anonymous):

is there a picture or anything ?

OpenStudy (brooke2353):

no its just a problem try to find the factor for each trinomial

OpenStudy (anonymous):

ok one minute

OpenStudy (brooke2353):

thank you i hafe trtied but i keep failing

OpenStudy (brooke2353):

have*

OpenStudy (anonymous):

3(4^2)-10*4=8 3*16-40=8 48-40=8 8=8

OpenStudy (anonymous):

(x + 4/3)(x + 2)

OpenStudy (ribhu):

3x^2+6x+4x+8 3x(x+2)+4(x+2) (x+2)(3x+4)

OpenStudy (ribhu):

factors

OpenStudy (brooke2353):

thank you ribhu